1. Briefly explain the following optical classification of materials:                  [3]

-Transparent materials


-Translucent materials


- Opaque materials




  1. A light beam in air hits a sheet of crown glass. Compute the velocity of light in crown glass.
  2. (nair = 1.0003, ncrown glass = 1.52, c = 3 × 108m/s).        
1
Expert's answer
2021-09-30T14:35:22-0400


TRANSPARENT -:

these are the materials which allows light to pass through them completely. and hence, we can see clearly through transparent objects.


TRANSLUCENT :-

these are the materials which allows light to pass through them but not completely. half of the light get absorbed. that's why we can't see through it clearly.


OPAQUE: -

these are the materials which don't allow light to pass through them. almost all light get absorbed. that's why we can't see through opaque objects.

For example


transparent - glass window


translucent - oily paper, frosted glass, glass used in bathroom.


opaque - wood, notebook, metal sheet.


"v=\frac{c}{n}"

"v_{air}=\frac{3\times10^{8}}{1.0003}=2.99\times10^{8}m/sec"


"v_{glass}=\frac{3\times10^8}{1.52}=1.97\times10^8m/sec"

"v_{air}>v_{glass}"